OpenStudy (anonymous):

Find the slope of the line through (3, –6) and (5, 8). (1 point) –7 7 –1over 7 1 over 7

OpenStudy (chriss):

\[\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (anonymous):

(8-(-6))/(5-3) = 14/2 = 7